OpenStudy (anonymous):

let f(x) = -2 + 4 and get g(x) = -6x - 7. find f(x) - g(x).

OpenStudy (the_fizicx99):

f(x) = -2 + 4 & g(x) = -6x - 7 Note: I think you're missing a x in -2 + 4 f(x) - g(x) = -2 + 4 - (-6x - 7) -2 + 4 + 6x - 7

OpenStudy (anonymous):

yes sorry its -2x + 4

OpenStudy (the_fizicx99):

Then its -2x + 4 + 6x -7

OpenStudy (anonymous):

so that is how you do functions, just add the two??

OpenStudy (the_fizicx99):

It's asking for f(x) - g(x) so yes, thats f(x) - g(x)

OpenStudy (anonymous):

Ok thank you

OpenStudy (the_fizicx99):

Yw